OpenStudy (anonymous):

FAN AND MEDAL!!! There is an inverse relationship between x and y. If y is twenty when x is 2, what is x when y is 8? 5 8 forty eighty

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

y = k / x xy = k = 40 x = 40 / y = 40 / 8 x = 5